OpenStudy (anonymous):

Piper invested $7,400 in a Certificate of Deposit $4,980 in a coperate bond, $5100 in a common stock and $6,350 in a preffered stock, The COD has a rate of return of 3.4%; the coperate bond is 5.8%; the common stock's rate of return is-1.9%; and the preffered stock has a rate of return of 4.6%. What is Piper's weighted mean rate of them?

OpenStudy (nurali):

7400x0.034 = 251.60 4980x0.058 = 288.84 5100x(-0.019) = -96.90 6350x0.046 = 735.64 Total invested = 7400+4980+5100+6350 = $23,830 Total return = 251.60+288.84-96.90+735.64 = $735.64 weighted mean rate of return = 735.64/23830 = .03087 = 3.09%